The question is , the lenght of a rectangle is twice the width. The area is 648sqrt. find the length and the width. I have never been very good at math and I have being trying to do this on my own but can not seem to understand how word problems work.
Answer by stanbon(75887)

The question is , the length of a rectangle is twice the width.
Draw the rectangle.
Label the width as x and the length as 2x.
Then the area = x*2x = 2x^2
-------------------------------------------------
Equation:
The area is 648sqrt. find the length and the width.
Area = length*width
Area = 2x*x = 2x^2
648 = 2x^2
324 = x^2
x = 18 (that is the width)
2x = 36 (that is the length)
.
-------------------------------
Cheers,
Stan H.